f(x)= 2x^-1 g(x)= x-2
f(g(x))

Answer by stanbon(75887)

f(x)= 2x^-1
g(x)= x-2
-----------------
f(g(x)) = f(x-2)
---
= 2(x-2)^(-1)
= 2/(x-2)
================
Cheers,
Stan H.
================